# Slice-Level Scheduling for High Throughput and Load Balanced LLM Serving

### v1

@@ 6. Related Works
Request scheduling. Existing serving systems (triton, ; tensorflowserving, ) lack model-specific optimizations and adopt the sequence-level scheduling to serve requests in an FCFS manner with a fixed batch size which leads to severe computational inefficiency for LLM serving. Orca (yu2022orca, ) and FastGen (deepspeed-fastgen, ) leverage the iteration-level scheduling to improve the serving efficiency but they adopt conservative memory management strategies which limit the throughput. On the shoulder of the aforementioned works, PiA (zheng2024response, ) and $S^{3}$ (jin2024s, ) propose to roughly predict the request generation length to enhance serving efficiency for sequence-level and iteration-level scheduling, respectively. However, they require significant computational overhead to fine-tune LLMs or train predictors while SCLS can support LLMs seamlessly without any additional computational overhead.
Low-level optimization. Kernel optimization (deepspeed-inference, ; dao2022flashattention, ; kwon2023efficient, ; dao2023flashattention, ) speeds up training and inference by implementing efficient CUDA kernels for transformer-based LLMs to improve computational efficiency. Parallelism optimization (huang2019gpipe, ; narayanan2019pipedream, ; narayanan2021efficient, ; miao2022galvatron, ; miao2023sdpipe, ) proposes to split the model at intra-layer and inter-layer granularity to enable distributed training and inference for LLMs. These studies are orthogonal to SCLS, and SCLS can be augmented with these low-level techniques.
DNN serving system. Recently proposed deep neural network (DNN) serving systems (romero2021infaas, ; gunasekaran2022cocktail, ; wang2023tabi, ) serve requests in a model-less manner, where developers simply specify the performance and accuracy requirements for their applications, and the serving system chooses a model for each query. Besides, (ali2020batch, ; ali2022optimizing, ; wu2022serverless, ) propose to leverage the automated resource management of serverless platforms to elastically handle requests. Moreover, Alpaserve (li2023alpaserve, ) improves the serving efficiency with statistical multiplexing. However, none of these systems are specifically optimized for generative LLMs.
7. Discussion
@@
Integration with continuous batching. For continuous batching, if the generation length of a request is too long, it will consume a large amount of memory, which causes newly arrived requests to queue up and can not be added to the serving process due to insufficient memory. By limiting the number of generated tokens for requests in each schedule to a small slice length, we can solve this problem by timely rescheduling the long request to an instance with more free memory. Besides, we can also accurately estimate the memory consumed by requests in each schedule to serve as many requests in parallel as possible without causing OOM errors and achieve balanced memory consumption across multiple LLM instances. We are working on implementing SCLS on top of vllm (vllm, ) to integrate with continuous batching, and we are also trying to mitigate the overhead of the prefill re-computation by proactively swapping the key-value cache between the CPU and GPU memory.
Fast support for inference engines. SCLS is a pure scheduling system. It can quickly support any inference engine that can serve requests with static batching without any modification of the engine. In this paper, we support huggingface-transformers and deepspeed-inference with only 350 lines of code, and we are making efforts to support more inference engines such as tensorrt-llm (tensorrt-llm, ).
Seamless support for LLMs. SCLS can be directly applied to all the transformer-based LLMs that generate tokens in an autoregressive manner, and SCLS does not require additional program development to support new models. Besides, no matter what engines the LLMs are loaded and run on, SCLS takes effect as long as those engines use the key-value cache mechanism during inference.
8. Conclusion
In this paper, we propose SCLS, a pure scheduling system to improve throughput for LLM serving and achieve load balancing across multiple LLM instances. We introduce the technical details of SCLS integration with static batching in this paper, where the number of iterations for each batch serving is limited to a small slice length. In this way, SCLS can obtain an accurate estimation of memory consumption and batch serving time, which enables effective scheduling. Since in actual scenarios the majority of requests have small generation lengths, the prefill re-computation overhead of SCLS is small when the slice length is set to an appropriate value. Extensive experiments are conducted to demonstrate that SCLS can far outperform existing ILS and SLS schedulers in terms of request throughput while greatly alleviating load imbalance. In the future, we will implement efficient CUDA kernels to achieve dedicated GPU memory management strategies for SCLS to further improve batch serving efficiency.

### v2

@@ VI-A Model Compression
Existing serving systems [2, 3] adopt the SLS to serve requests in an FCFS manner with a fixed batch size. Orca [4] and Deepspeed-FastGen [6] leverage the ILS to improve the serving efficiency, but they adopt conservative memory management strategies that limit throughput. PiA [16] and $S^{3}$ [17] propose to roughly predict the request generation length to enhance serving efficiency. However, they require significant computational overhead to fine-tune LLMs or train predictors. Recent studies [37, 38, 39, 40, 41, 42, 20, 43, 18, 44, 14, 45] propose a series of advanced techniques such as chunked prefill, disaggregated inference, and multi-lora serving. These works are designed for continuous batching, and we will combine these techniques with generation slicing in the future.
VII Generation Slicing for Continuous Batching
@@
For continuous batching, if the generation length of a request is too long, it will consume a large amount of memory, which causes newly arrived requests to queue up and can not be added to the serving process due to insufficient memory. By limiting the number of generated tokens for requests in each schedule to a small slice length, we can solve this problem by timely re-scheduling the long request to an instance with more free memory. Besides, we can also accurately estimate the memory consumed by requests in each schedule to serve as many requests in parallel as possible without causing OOM errors or request preemption and achieve balanced memory consumption across multiple LLM instances. Besides, for requests re-scheduled to the same LLM instance, we can mitigate the prefill re-computation overhead by reusing the previously produced key-value cache with prefix-caching [46] or proactively swapping the key-value cache between GPU and CPU memories. For requests re-scheduled to different LLM instances, the prefill re-computation can be amortized with decoding using chunked prefill [20].
VIII Conclusion and Future Works
In this paper, we propose the core idea of generation slicing and design slice-level scheduling (SCLS) to achieve high-throughput and load-balanced LLM serving with static batching. By splitting the predefined maximal generation length limit into slices, SCLS can provide a precise range of serving time and memory usage for batches, laying the foundation for the design of adaptive bathing and max-min offloading algorithms. Extensive experiments are conducted to confirm the superiority of SCLS in throughput improvement and load balance. In the future, we will apply generation slicing to continuous batching and mitigate the overhead of the prefill re-computation by amortizing prefill computation using the chunked prefill technique and proactively swapping the key-value cache between the CPU and GPU memory.